# RewardMap: Tackling Sparse Rewards in Fine-grained Visual Reasoning via Multi-Stage Reinforcement Learning

### v1

@@ 1 Introduction
Fine-grained visual reasoning over structured visual inputs remains a significant challenge for multimodal large language models (MLLMs) (Bai et al., 2025; OpenAI, 2025; Zhang et al., 2025b). Recently, ReasonMap (Feng et al., 2025b) was introduced as a benchmark on high-resolution transit maps in the real world, where tasks (e.g., route planning) combine visual understanding with spatial reasoning, jointly constituting the fine-grained visual reasoning challenge. This task is not only of practical value for real-world navigation and transportation systems, but also of fundamental scientific interest as it exposes reasoning gaps in current MLLMs. Despite steady advances in vision–language pre-training (Liu et al., 2023; 2024; Bai et al., 2025), existing models consistently struggle with the visual and spatial reasoning demands in ReasonMap. This gap motivates us to investigate how reinforcement learning (RL) (Zhang et al., 2025a; Shao et al., 2024; Guo et al., 2025a) can be adapted to enhance fine-grained visual reasoning abilities in structured visual domains such as transit maps.
@@
However, directly applying standard RL methods (Rafailov et al., 2023; Shao et al., 2024) to complex tasks such as ReasonMap is highly challenging, as supervision signals are inherently sparse (i.e., success is typically judged only at the final answer after a long reasoning chain (Quadros et al., 2025; Cao et al., 2024)). The difficulty of the tasks further amplifies this sparsity, which in turn destabilizes optimization and hinders effective exploration (Wang et al., 2025a). While classical approaches like Supervised Fine-Tuning (SFT) (Liu et al., 2023; Wei et al., 2025) offer dense supervision, they fall short in equipping models for the long-chain decision-making intrinsic to visual reasoning tasks. This mismatch between task complexity and supervision signal forms a critical bottleneck in leveraging RL for fine-grained visual reasoning.
@@
To address this issue, we first construct ReasonMap-Plus, an extended dataset that introduces dense reward signals for further cold-start training. Tasks in ReasonMap-Plus are organized along a natural difficulty continuum, from simple Visual Question Answering (VQA) that enhances perception to progressively harder visual tasks reflecting the complexity of fine-grained visual reasoning queries. We further introduce RewardMap, a multi-stage RL framework with detailed reward design. It consists of two key components: (1) a reward scheme that, beyond basic format and correctness rewards, incorporates detail rewards to mitigate sparsity in supervision for hard samples and adopts a difficulty-aware design to account for task complexity; and (2) a cold-start strategy that departs from SFT-based initialization (Xu et al., 2024; Guo et al., 2024; Wei et al., 2025) by directly employing RL, ensuring alignment between reward signals and task objectives from the outset. Training data are organized from easy to hard across multiple stages, with dense and accessible rewards at lower levels supporting effective cold-start training. This staged strategy systematically bridges perception and reasoning within a unified RL framework.
@@
We conduct extensive experiments on ReasonMap and ReasonMap-Plus to evaluate the effectiveness of RewardMap. Results indicate that each component yields measurable gains, with their integration delivering the best overall performance. Moreover, beyond the targeted benchmarks, models trained with RewardMap achieve consistent improvements ($3.47\%$ average) across six benchmarks  (Wu & Xie, 2024; Wang et al., 2024; Li et al., 2024; Wang et al., 2025d; Masry et al., 2022; Chen et al., 2024a) covering spatial reasoning, fine-grained visual reasoning, and general tasks, suggesting enhanced general visual perception and reasoning capabilities.
@@ 2 Related Work
Visual Reasoning in MLLMs. The development of MLLMs has rapidly progressed from foundational models that bridge vision and language encoders, such as Flamingo (Alayrac et al., 2022), to those enhanced by visual instruction tuning like LLaVA (Liu et al., 2023; 2024). To elicit more complex, step-by-step reasoning, researchers have adapted Chain-of-Thought (CoT) (Wei et al., 2022) prompting from the language domain to the multimodal context (MCoT) (Zhang et al., 2023). However, a key limitation of early MCoT is its reliance on purely textual rationales (Wang et al., 2025e), which can be a bottleneck for expressing fine-grained visual logic (Zhang et al., 2025b). More recent work has thus focused on developing vision-centric reasoning processes (Dong et al., 2025; Man et al., 2025), such as generating intermediate visual representations as perception tokens to aid the reasoning chain (Bigverdi et al., 2025). Despite these advances, a significant performance gap persists. Benchmarks specifically designed to test abstract, spatial, and logical reasoning, such as VisuLogic (Xu et al., 2025) and ReasonMap (Feng et al., 2025b), reveal that even state-of-the-art models struggle with tasks that require high-fidelity visual and topological understanding, underscoring the need for new methods tailored for structured visual domains, such as transit maps.
Reinforcement Learning for Visual Reasoning. Reinforcement Learning (RL) (Wu et al., 2025; Sarch et al., 2025; Feng et al., 2025a) offers a powerful paradigm for enhancing the reasoning capabilities of MLLMs beyond the static nature of Supervised Fine-Tuning (SFT) (Liu et al., 2023; Wei et al., 2025). The application of RL in this domain has evolved from methods like Reinforcement Learning from Human Feedback (RLHF) (Bai et al., 2022) to more direct and stable policy optimization algorithms such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024). While SFT on CoT datasets can activate a model’s latent reasoning abilities (Xu et al., 2024; Guo et al., 2024; Wei et al., 2025), it often leads to overfitting and cognitive rigidity, limiting generalization (Chu et al., 2025; Shen et al., 2025; Wang et al., 2025b). RL provides a dynamic alternative by enabling models to learn a reasoning policy through exploration and reward (Guo et al., 2025a; Zhang et al., 2025a). However, its direct application to complex visual reasoning is often hindered by sparse rewards (Quadros et al., 2025; Cao et al., 2024), where a signal is only provided upon completion of a long reasoning chain (Wang et al., 2025a). This motivates us to develop a framework that can effectively apply RL to structured visual tasks by mitigating sparse reward problems from the outset.
@@ 5.1 Experimental Setups
Training Details. We conduct training experiments under various settings with the training data from ReasonMap ($696$ samples, $R_{train}$(Feng et al., 2025b) and our proposed ReasonMap-Plus ($2,570$ samples, $RPlus_{train}$) on Qwen2.5-VL models (Bai et al., 2025). For GRPO (Shao et al., 2024) RL training, we use AdamW with an initial learning rate of $1.0\times 10^{-6}$ and a KL divergence coefficient of $1.0\times 10^{-3}$. The global batch size is $16$, and we sample $8$ responses per query. Besides RL training, we conduct baseline experiments using SFT with training data from ReasonMap-Plus. We apply LoRA (Hu et al., 2022) to the language blocks and the text-vision projector with an initial learning rate of $1.0\times 10^{-4}$. For implementation, we adopt LLaMA-Factory (Zheng et al., 2024) and VeRL (Sheng et al., 2024). All experiments were conducted on $8$ NVIDIA H800 GPUs.
Inference Details. All evaluations use greedy decoding (temperature=0). For open-source models, we cap the maximum output length at $2,048$ tokens and retain all other settings from the official HuggingFace configurations; deployments use PyTorch with Transformers [^3] https://github.com/huggingface/transformers on $8$ NVIDIA H800 GPUs. For closed-source models, we evaluate through official APIs using default settings. The evaluated models include: Kimi-VL-A3B-Insturct/Thinking (Team et al., 2025), Qwen2.5-VL-3/7/32/72B-Instruct (Bai et al., 2025), and Seed1.5-VL (Guo et al., 2025b).
Evaluation Datasets. We first evaluate on the test sets of ReasonMap and ReasonMap-Plus, and we report the metrics adjusted by the difficulty-aware weighting scheme (see details in Appendix C.1). To further assess the capability gains brought by RewardMap, we next employ six widely-used benchmarks spanning three dimensions (e.g., spatial reasoning, fine-grained visual reasoning, and general tasks): SEED-Bench-2-Plus (Li et al., 2024), SpatialEval (Wang et al., 2024), VBench Wu & Xie (2024), HRBench (Wang et al., 2025d), ChartQA (Masry et al., 2022), and MMStar (Chen et al., 2024a) (see Appendix C.2 for details). All evaluations are conducted with VLMEvalKit [^4] https://github.com/open-compass/VLMEvalKit.
5.2 Main Results
@@
+ RL (baseline) $R_{train}$ $26.22\%$ / $26.04\%$ $5.52$ / $9.52$ $44.64\%$ $37.57\%$ / $52.01\%$
@@
Results on ReasonMap and ReasonMap-Plus We evaluate the proposed RewardMap on ReasonMap and ReasonMap-Plus, both of which provide fine-grained difficulty annotations to assess visual understanding, visual reasoning, and spatial reasoning. In addition to the baselines introduced in Section 4.1, we compare against two widely-used settings: (1) SFT $\rightarrow$ RL baseline, which applies SFT on ReasonMap-Plus followed by RL on ReasonMap, and (2) RL baseline, which performs RL on the combined training data from ReasonMap and ReasonMap-Plus. For both baselines, the reward design includes only format and correctness terms.
@@
Results on Other Benchmarks We further evaluate the generalization ability of our method on six benchmarks spanning three task categories introduced in Section 5.1. Table 2 reports the results of reference models, the SFT $\rightarrow$ RL baseline, and our proposed RewardMap. Across all six benchmarks, RewardMap achieves consistent improvements, with the most substantial gain of $13.51\%$ observed on SpatialEval. While the SFT $\rightarrow$ RL baseline also yields stable improvements, its performance remains inferior to RewardMap. These results highlight the significant contributions of both RewardMap and the datasets (both ReasonMap and ReasonMap-Plus) to enhancing model general capability.
@@ 5.3 Qualitative Results
Figure 3 presents a qualitative comparison between reference models, the baseline RL model, and our proposed RewardMap. Across diverse maps, we observe that reference models and the baseline RL model often suffer from visual confusion (e.g., mistaking the route or stop as shown by Example a & c in Figure 3) or even hallucination (e.g., repeating the same route many times in Figure 3 (b)). In contrast, RewardMap consistently produces the correct target routes. These examples highlight the effectiveness of RewardMap in handling visually complex maps (e.g., Example a & b in Figure 3) and reducing both visual confusions and hallucinations.
@@ 5.4 Diagnostic Experiments
We provide extensive ablation studies with Qwen2.5-VL-7B-Instruct in this section, along with verification of RewardMap tackling sparse reward and across different model scales (e.g., Qwen2.5-VL-3B-Instruct).
@@
Effectiveness of RewardMap across Model Scales. In the end, we evaluate the effectiveness of RewardMap across different model scales. Due to training cost constraints, we adopt Qwen2.5-VL-3B-Instruct as the base model and compare the baseline RL with RewardMap. As shown in Table 6, RewardMap achieves the promising results, demonstrating its robustness and effectiveness.

### v2

@@ 1 Introduction
Fine-grained visual reasoning over structured visual inputs remains a significant challenge for multimodal large language models (MLLMs) (Bai et al., 2025b; OpenAI, 2025; Zhang et al., 2025c; Liang et al., 2025; Li et al., 2025c; d). Recently, ReasonMap (Feng et al., 2025b) was introduced as a benchmark on high-resolution transit maps in the real world, where tasks (e.g., route planning) combine visual understanding with spatial reasoning, jointly constituting the fine-grained visual reasoning challenge (Xie et al., 2025a; Hu et al., 2025b; Kong et al., 2025b; a). This task is not only of practical value for real-world navigation and transportation systems, but also of fundamental scientific interest as it exposes reasoning gaps in current MLLMs. Despite steady advances in vision–language pre-training (Liu et al., 2023; 2024; Bai et al., 2025b), existing models consistently struggle with the visual and spatial reasoning demands in ReasonMap. This gap motivates us to investigate how reinforcement learning (RL) (Zhang et al., 2025a; Shao et al., 2024; Guo et al., 2025a) can be adapted to enhance fine-grained visual reasoning abilities in structured visual domains such as transit maps.
@@
However, directly applying standard RL methods (Rafailov et al., 2023; Shao et al., 2024) to complex tasks such as ReasonMap is highly challenging, as supervision signals are inherently sparse, i.e., success is typically judged only at the final answer after a long reasoning chain (Quadros et al., 2025; Cao et al., 2024; Chen et al., 2025). The difficulty of the tasks further amplifies this sparsity, which in turn destabilizes optimization and hinders effective exploration (Wang et al., 2025a). While classical approaches like Supervised Fine-Tuning (SFT) (Liu et al., 2023; Wei et al., 2025) offer dense supervision, they fall short in equipping models for the long-chain decision-making intrinsic to visual reasoning tasks. This mismatch between task complexity and supervision signal forms a critical bottleneck in leveraging RL for fine-grained visual reasoning.
@@
To address this issue, we first construct ReasonMap-Plus, an extended dataset that introduces dense reward signals for further cold-start training. Tasks in ReasonMap-Plus are organized along a natural difficulty continuum, from simple Visual Question Answering (VQA) that enhances perception to progressively harder visual tasks reflecting the complexity of fine-grained visual reasoning queries. We further introduce RewardMap, a multi-stage RL framework with detailed reward design. It consists of two key components: (1) a reward scheme that, beyond basic format and correctness rewards, incorporates detail rewards to mitigate sparsity in supervision for hard samples and adopts a difficulty-aware design to account for task complexity; and (2) a cold-start strategy that departs from SFT-based initialization (Xu et al., 2024; Guo et al., 2024; Wei et al., 2025; Yan et al., 2025) by directly employing RL, ensuring alignment between reward signals and task objectives from the outset. Training data are organized from easy to hard across multiple stages, with dense and accessible rewards at lower levels supporting effective cold-start training. This staged strategy systematically bridges perception and reasoning within a unified RL framework.
@@
We conduct extensive experiments on ReasonMap and ReasonMap-Plus to evaluate the effectiveness of RewardMap. Results indicate that each component yields measurable gains, with their integration delivering the best overall performance. Moreover, beyond the targeted benchmarks, models trained with RewardMap achieve consistent improvements ($3.47\%$ average) across six benchmarks  (Wu & Xie, 2024; Wang et al., 2024a; Li et al., 2024; Wang et al., 2025f; Masry et al., 2022; Chen et al., 2024a) covering spatial reasoning, fine-grained visual reasoning, and general tasks, suggesting enhanced general visual perception and reasoning capabilities.
@@ 2 Related Work
Visual Reasoning in MLLMs. The development of MLLMs has rapidly progressed from foundational models that bridge vision and language encoders, such as Flamingo (Alayrac et al., 2022), to those enhanced by visual instruction tuning like LLaVA (Liu et al., 2023; 2024). To elicit more complex, step-by-step reasoning, researchers have adapted Chain-of-Thought (CoT) (Wei et al., 2022) prompting from the language domain to the multimodal context (MCoT) (Zhang et al., 2023). However, a key limitation of early MCoT is its reliance on purely textual rationales (Wang et al., 2025g), which can be a bottleneck for expressing fine-grained visual logic (Zhang et al., 2025c). More recent work has thus focused on developing vision-centric reasoning processes (Dong et al., 2025; Man et al., 2025), such as generating intermediate visual representations as perception tokens to aid the reasoning chain (Bigverdi et al., 2025). Despite these advances, a significant performance gap persists. Benchmarks specifically designed to test abstract, spatial, and logical reasoning, such as VisuLogic (Xu et al., 2025) and ReasonMap (Feng et al., 2025b), reveal that even state-of-the-art models struggle with tasks that require high-fidelity visual and topological understanding, underscoring the need for new methods tailored for structured visual domains, such as transit maps.
@@
Reinforcement Learning for Reasoning. Reinforcement Learning (RL) (Wu et al., 2025a; Sarch et al., 2025; Feng et al., 2025a) has emerged as a powerful paradigm for improving the reasoning capabilities of multimodal and language-only models beyond the static nature of Supervised Fine-Tuning (SFT) (Liu et al., 2023; Wei et al., 2025). The evolution of RL for reasoning spans from traditional RLHF pipelines (Bai et al., 2022) to more stable and direct policy optimization objectives such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024). Recent work further demonstrates the effectiveness of structured or curriculum-based RL strategies for enhancing LLM reasoning. Curriculum RL (Parashar et al., 2025) shows that progressing from easy to hard tasks can significantly strengthen reasoning robustness; Kimi K1.5 (Team et al., 2025) adopts multi-stage RL to improve long-chain and tool-integrated reasoning; and Logic-RL (Xie et al., 2025b) leverages logical structure to guide policy optimization. While SFT on CoT-rich datasets can activate latent reasoning abilities (Xu et al., 2024; Guo et al., 2024; Wei et al., 2025), it often leads to overfitting and cognitive rigidity (Chu et al., 2025; Shen et al., 2025; Wang et al., 2025b). RL provides a dynamic alternative by allowing models to learn a reasoning policy through exploration and reward (Guo et al., 2025a; Zhang et al., 2025a). However, its direct application to complex visual reasoning remains challenging due to sparse reward signals (Quadros et al., 2025; Cao et al., 2024), where supervision is only provided at the end of long reasoning trajectories (Wang et al., 2025a). These limitations motivate the development of an RL framework tailored for structured visual tasks that alleviates sparse rewards from the outset.
@@ 5.1 Experimental Setups
+ RL (GRPO) $R_{train}$ $26.22\%$ / $26.04\%$ $5.52$ / $9.52$ $44.64\%$ $37.57\%$ / $52.01\%$
+ RL (REINFORCE++) $R_{train}$ $27.17\%$ / $27.60\%$ $5.68$ / $10.12$ $44.64\%$ $36.82\%$ / $52.79\%$
+ RL (ReMax) $R_{train}$ $26.22\%$ / $27.26\%$ $5.57$ / $9.99$ $45.39\%$ $38.37\%$ / $52.70\%$
@@
Training Details. We conduct training experiments under various settings with the training data from ReasonMap ($696$ samples, $R_{train}$(Feng et al., 2025b) and our proposed ReasonMap-Plus ($2,570$ samples, $RPlus_{train}$) on Qwen2.5-VL models (Bai et al., 2025b) with $8$ NVIDIA H800 GPUs. For GRPO (Shao et al., 2024) RL training, we use AdamW with an initial learning rate of $1.0\times 10^{-6}$ and a KL divergence coefficient of $1.0\times 10^{-3}$. The global batch size is $16$, and we sample $8$ responses per query. Besides RL training, we conduct baseline experiments using SFT with training data from ReasonMap-Plus. We apply LoRA (Hu et al., 2022) to the language blocks and the text-vision projector with an initial learning rate of $1.0\times 10^{-4}$. For implementation, we adopt LLaMA-Factory (Zheng et al., 2024) and VeRL (Sheng et al., 2024). Additionally, we include two RL training baselines with REINFORCE++ (Hu et al., 2025a) and ReMax (Li et al., 2023).
Inference Details. All evaluations use greedy decoding (temperature=0). For open-source models, we cap the maximum output length at $2,048$ tokens and retain all other settings from the official HuggingFace configurations; deployments use PyTorch with Transformers [^3] https://github.com/huggingface/transformers on $8$ NVIDIA H800 GPUs. For closed-source models, we evaluate through official APIs using default settings. The evaluated models include: Kimi-VL-A3B-Insturct/Thinking (Team et al., 2025), Qwen2.5-VL-3/7/32/72B-Instruct (Bai et al., 2025b), Seed1.5-VL (Guo et al., 2025b), GPT-5/4o (OpenAI, 2024).
Evaluation Datasets. We first evaluate on the test sets of ReasonMap and ReasonMap-Plus, and we report the metrics adjusted by the difficulty-aware weighting scheme (see details in Appendix C.1). To further assess the capability gains brought by RewardMap, we next employ six widely-used benchmarks spanning three dimensions (e.g., spatial reasoning, fine-grained visual reasoning, and general tasks): SEED-Bench-2-Plus (Li et al., 2024), SpatialEval (Wang et al., 2024a), VBench Wu & Xie (2024), HRBench (Wang et al., 2025f), ChartQA (Masry et al., 2022), and MMStar (Chen et al., 2024a) (see Appendix C.2 for details). All evaluations are conducted with VLMEvalKit [^4] https://github.com/open-compass/VLMEvalKit.
5.2 Main Results
Results on ReasonMap and ReasonMap-Plus. We evaluate the proposed RewardMap on ReasonMap and ReasonMap-Plus, both of which provide fine-grained difficulty annotations to assess visual understanding, visual reasoning, and spatial reasoning. In addition to the baselines introduced in Section 4.1, we compare against two widely-used settings: (1) SFT $\rightarrow$ RL baseline, which applies SFT on ReasonMap-Plus followed by RL on ReasonMap, and (2) RL baseline, which performs RL on the combined training data from ReasonMap and ReasonMap-Plus. For both baselines, the reward design includes only format and correctness terms.
@@ 5.2 Main Results
Results on Other Benchmarks. We further evaluate the generalization ability of our method on six benchmarks spanning three task categories introduced in Section 5.1. Table 2 reports the results of reference models, the SFT $\rightarrow$ RL baseline, and our proposed RewardMap. Across all six benchmarks, RewardMap achieves consistent improvements, with the most substantial gain of $13.51\%$ observed on SpatialEval. While the SFT $\rightarrow$ RL baseline also yields stable improvements, its performance remains inferior to RewardMap. These results highlight the significant contributions of both RewardMap and the datasets (both ReasonMap and ReasonMap-Plus) to enhancing model general capability. Beyond the aggregate results, we further report task-wise improvements. On SEED-Bench-2-Plus, RewardMap improves the map split. For SpatialEval, accuracy rises substantially on mazenav ($19.60\%\rightarrow 57.20\%$) and moderately on spatialreal ($70.37\%\rightarrow 72.59\%$). On HRBench, performance increases for both single-image reasoning ($85.25\%\rightarrow 88.00\%$) and cross-image alignment ($52.25\%\rightarrow 54.00\%$). For MMStar, RewardMap achieves steady gains across fine-grained perception ($59.60\%\rightarrow 60.80\%$), instance reasoning ($68.00\%\rightarrow 71.20\%$), and mathematical reasoning ($59.60\%\rightarrow 63.20\%$). These results show that RewardMap consistently strengthens visual grounding and robustness in diverse spatial and reasoning tasks.
@@ 5.3 Qualitative Results
Figure 3 presents a qualitative comparison between reference models, the baseline RL model, and our proposed RewardMap. Across diverse maps, we observe that reference models and the baseline RL model often suffer from visual confusion (e.g., mistaking the route or stop as shown by Example a & c in Figure 3) or even hallucination (e.g., repeating the same route many times in Figure 3 (b)). In contrast, RewardMap consistently produces the correct target routes. These examples highlight the effectiveness of RewardMap in handling visually complex maps (e.g., Example a & b in Figure 3) and reducing both visual confusions and hallucinations. Additionally, we present further comparison cases that provide additional evidence that RewardMap improves visual grounding and reduces the likelihood of visual confusion or hallucination. These examples are included in Appendix D.1.
@@
Table 6: Ablation on the hyperparameter $\alpha$. “S.S.” represents results for short questions, while “L.L.” denotes results for long questions.
Model Training Data ReasonMap ($S./L.$) ReasonMap-Plus
+ RewardMap $\alpha$=0.3] $RPlus_{train}$ + $R_{train}$ $30.73\%$ / $31.16\%$ $6.01$ / $10.81$ $72.08\%$ $68.41\%$ / $75.91\%$
+ RewardMap $\alpha$=0.5] $RPlus_{train}$ + $R_{train}$ $31.51\%$ / $31.77\%$ $6.21$ / $11.22$ $74.25\%$ $72.18\%$ / $76.42\%$
+ RewardMap $\alpha$=0.7] $RPlus_{train}$ + $R_{train}$ $32.03\%$ / $32.20\%$ $6.23$ / $11.20$ $72.81\%$ $70.14\%$ / $75.59\%$
Table 7: Ablation on the difficulty-aware weights. “S.S.” represents results for short questions, while “L.L.” denotes results for long questions.
Model Training Data ReasonMap ($S./L.$) ReasonMap-Plus
+ Reward Design $\gamma_{e/m/h}=(1.0,1.1,1.2)$] $R_{train}$ $26.91\%$ / $28.47\%$ $5.64$ / $10.24$ $45.34\%$ $37.97\%$ / $53.02\%$
+ Reward Design $\gamma_{e/m/h}=(1.0,1.2,1.5)$] $R_{train}$ $29.08\%$ / $29.95\%$ $5.88$ / $10.53$ $45.16\%$ $37.79\%$ / $52.84\%$
+ Reward Design $\gamma_{e/m/h}=(1.0,1.5,2.0)$] $R_{train}$ $6\% / 28.9928$$9\%$ $5.93$ / $10.31$ $45.16\%$ $37.31\%$ / $53.35\%$
+ Reward Design $\beta_{0/1}=(0.0,0.2)$] $R_{train}$ $28.30\%$ / $29.34\%$ $5.68$ / $10.12$ $44.60\%$ $37.39\%$ / $52.10\%$
+ Reward Design $\beta_{0/1}=(0.0,0.5)$] $R_{train}$ $29.08\%$ / $29.95\%$ $5.88$ / $10.53$ $45.16\%$ $37.79\%$ / $52.84\%$
+ Reward Design $\beta_{0/1}=(0.0,0.8)$] $R_{train}$ $29.60\%$ / $29.34\%$ $5.89$ / $10.23$ $45.39\%$ $38.68\%$ / $52.38\%$
@@ 5.4 Diagnostic Experiments
We provide extensive ablation studies using Qwen2.5-VL-7B-Instruct, and further verify the effectiveness of RewardMap in addressing sparse-reward issues, across different model scales (e.g., Qwen2.5-VL-3B-Instruct) and different model architectures (e.g., Kimi-VL-A3B-Instruct).
@@
Ablation on the Hyperparameter $\alpha$. Based on the full RewardMap pipeline, we varied only $\alpha$. The results are reported in Table 6. We observe a slight performance drop when $\alpha$ is small, while the performance becomes similar once $\alpha$ enters a higher range. These results suggest that the detailed reward is indeed beneficial (as $\alpha$ determines its contribution to the total reward), and they also indicate that $\alpha$ is not sensitive within a reasonable interval.
@@
Ablation on the Difficulty-Aware Weighting Scheme. Using the ReasonMap training set, we further conduct two ablations: one varying only the relative magnitudes of $\gamma_{e/m/h}$ and another varying only those of $\beta_{0/1}$. As shown in Table 7, performance decreases slightly when the weights become nearly uniform, but remains stable once clear difficulty distinctions are introduced. These results support the effectiveness of our difficulty-aware weighting—where $\gamma_{e/m/h}$ and $\beta_{0/1}$ modulate map and question difficulty, respectively—and indicate that both weights are insensitive within a reasonable range.
@@
Effectiveness of RewardMap across Model Scales. We evaluate the effectiveness of RewardMap across different model scales. Due to training cost constraints, we adopt Qwen2.5-VL-3B-Instruct as the base model and compare the baseline RL with RewardMap. As shown in Table 8, RewardMap achieves the promising results, demonstrating its robustness and effectiveness.
Table 9: Results on Kimi-VL (Kimi-VL-A3B-Instruct). “S.S.” represents results for short questions, while “L.L.” denotes results for long questions.
Model Training Data ReasonMap ($S./L.$) ReasonMap-Plus
+ RewardMap $R_{train}$ $18.58\%$ / $17.36\%$ $4.70$ / $7.69$ $35.92\%$ $15.20\%$ / $57.50\%$
Generalization of RewardMap across Model Architectures. We further trained Kimi-VL (Kimi-VL-A3B-Instruct) with the RewardMap pipeline using the training set from ReasonMap. As shown in Table 9, the model achieves substantial performance gains, demonstrating that our method generalizes beyond the Qwen2.5-VL series.